# UniAudio: An Audio Foundation Model Toward Universal Audio Generation

### v5

@@ Abstract
Large Language models (LLM) have demonstrated the capability to handle a variety of generative tasks. This paper presents the UniAudio system, which, unlike prior task-specific approaches, leverages LLM techniques to generate multiple types of audio (including speech, sounds, music, and singing) with given input conditions. UniAudio 1) first tokenizes all types of target audio along with other condition modalities, 2) concatenates source-target pair as a single sequence, and 3) performs next-token prediction using LLM. Also, a multi-scale Transformer model is proposed to handle the overly long sequences caused by the residual vector quantization-based neural codec in tokenization. Training of UniAudio is scaled up to 165K hours of audio and 1B parameters, based on all generative tasks, aiming to obtain sufficient prior knowledge not only in the intrinsic properties of audio but also the inter-relationship between audio and other modalities. Therefore, the trained UniAudio model has the potential to become a foundation model for universal audio generation: it shows strong capability in all trained tasks and can seamlessly support new audio generation tasks after simple fine-tuning. Experiments demonstrate that UniAudio achieves state-of-the-art or at least competitive results on most of the 11 audio generation tasks. Demo and code are released. [^1] https://uniaudio666.github.io/demo_UniAudio/ * Equal contribution; $\dagger$ Corresponding author.
@@ 1 Introduction
Audio generation is an important component of generative AI. Recently, the popularity of generative AI has induced increasingly emergent and varying needs in audio generation: audio is expected to be generated based on humans’s demands, such as speech synthesis (TTS), voice conversion (VC), singing voice synthesis (SVS), text-to-sound, and text-to-music. Prior works on audio generation tasks are commonly task-specific: their designs heavily leverage domain knowledge and their usage is restricted to fixed setups (Tan et al., 2021; Luo & Mesgarani, 2019; Zmolikova et al., 2023; Huang et al., 2021b; Cho et al., 2021). Instead of taking care of each task independently, this work is an attempt to achieve universal audio generation, which intends to accomplish multiple audio generation tasks with only one unified model. The universal audio generation model is expected to obtain sufficient prior knowledge in audio and related modalities, which has the potential to provide simple and effective solutions for the increasing needs of generating diverse types of audio.
The superiority of Large Languge Models (LLM) in text-generative tasks inspires a series of LLM-based models in audio generation (Wang et al., 2023a; Kharitonov et al., 2023; Huang et al., 2023b; Agostinelli et al., 2023; Borsos et al., 2023). Among these works, LLM’s capability in independent tasks has been extensively studied in tasks like text-to-speech (TTS) (Wang et al., 2023a; Kharitonov et al., 2023; Huang et al., 2023b) and music generation (Agostinelli et al., 2023; Copet et al., 2023), and achieves competitive performance. However, LLM’s ability to process multiple tasks with a unified model is less exploited in audio generation research: most existing LLM-based works are still designed for single tasks (Wang et al., 2023a; Kharitonov et al., 2023). We argue that achieving universality and versatility in audio generation through the LLM paradigm is promising but has not yet been comprehensively studied before this work.
Toward universal audio generation, this work presents UniAudio, which adopts LLM techniques and is able to generate multiple types of audio (speech, sounds, music, and singing) conditioned on various input modalities, such as phoneme sequences, textual descriptions, and audio itself. The proposed UniAudio is mainly featured as follows: First, all types of audio, along with all other input modalities, are tokenized as discrete sequences. Specifically, a universal neural codec model is built to effectively tokenize audio regardless of the audio type, and other tokenizers are used to tokenize other different modalites. Then, UniAudio concatenates the source-target pair as a single sequence. Lastly, UniAudio performs next-token prediction using LLM. The residual vector quantization (Zeghidour et al., 2021) based on neural codecs is used in the tokenization process, resulting in overly long token sequences (one frame corresponding to multiple tokens) that cannot be processed efficiently by LLM. A multi-scale Transformer architecture is designed to reduce computational complexity by modeling the inter- and intra-frame correlation separately. Specifically, a global Transformer module is used to model the inter-frame correlation (e.g. semantic level), and a local Transformer module is used to model the intra-frame correlation (e.g. acoustic level).
To demonstrate the scalability of UniAudio for new tasks, the building process of UniAudio takes two stages. Firstly, the proposed UniAudio is trained on multiple audio generation tasks jointly, which allows the model to obtain sufficient prior knowledge not only of the intrinsic properties of audio but also of the interrelationship between audio and other input modalities. Secondly, through fine-tuning, the trained model can seamlessly support more unseen audio generation tasks. Thus, UniAudio has the potential to become a foundation model for universal audio generation: it is able to continuously support emergent needs in audio generation. Experimentally, our UniAudio supports 11 audio generation tasks: the training stage includes 7 audio generation tasks, while 4 tasks are further added in the fine-tuning stage. The building process of UniAudio is scaled up to 165k hours of audio and 1B parameters. Among the 11 tasks, UniAudio consistently obtains competitive performance in both objective and subjective evaluations. State-of-the-art results are even achieved on most of these tasks. Further investigation suggests that training multiple tasks simultaneously in the training stage is mutually beneficial to each task involved. In addition, UniAudio can effectively adapt to new audio generation tasks and outperform task-specific models with a non-trivial gap.
To sum up, this work reveals that building universal audio generation models is necessary, promising, and beneficial. The main contributions of this work are summarized as follows:
(1) Toward universal audio generation, UniAudio is presented as a unified solution for 11 audio generation tasks.
(2) Per methodology, UniAudio provides novel approaches for (i) sequential representations of audio and other input modalities; (ii) uniform formulation for LLM-based audio generation tasks; and (iii) efficient model architecture specifically designed for audio generation.
(3) Per experiments, the overall performance of UniAudio is well validated, and the benefits of building a versatile audio generation model are verified by exhaustive experimental results.
(4) Demo and code are released, in the hope that UniAudio can become a foundation model that supports emergent audio generation in future research.
2 UniAudio
This section introduces the technical details of the proposed UniAudio. Section 2.1 explains how audio and other modalities are tokenized. Then, all considered audio generation tasks are uniformly formulated in Section 2.2. Subsequently, the multi-scale Transformer architecture is proposed in Section 2.3 to handle the overly long sequence challenge caused by the adoption of neural codecs.
2.1 Tokenization
LLM are commonly used for sequential modeling, so audio and all other input modalities are tokenized before being processed. These processes for each modality are completed by independent modules. All of these modules are fixed in the optimization of UniAudio or parameter-free.
2.1.1 Audio
For all audio generation tasks considered in this work, audio, regardless of its types (speech, sounds, music, or singing), is the target to predict. Instead of modeling different types of audio separately, UniAudio intends to tokenize all types of audio as a single and unified modality (even though they commonly have distinct patterns, such as frequency span), which requires a model that is well-suited to mapping all audio types into a shared latent space. Following Wang et al. (2023a); Kharitonov et al. (2023), neural codec models (Défossez et al., 2022; Yang et al., 2023b; Kumar et al., 2023) are used in this work for audio tokenization. An audio signal of duration $d$ with sample rate $f_{s}$ can be represented by a sequence $\mathbf{x}\in[-1,1]^{d*f_{s}}$. An audio neural codec intends to compress ${\mathbf{x}}$ and then recover it as $\mathbf{\hat{x}}$ using an encoder-decoder architecture with a quantization module:
$\mathbf{h}=\text{Encoder}(\mathbf{x})\in\mathcal{R}^{T*L};\quad\mathbf{\hat{h}}=\text{Quantization}(\mathbf{h});\quad\mathbf{\hat{x}}=\text{Decoder}(\mathbf{\hat{h}})$ (1)
where $T$ denotes the number of audio frames after down-sampling in the encoder, and $L$ denotes the feature dimension of the encoder. The discrete representations of audio are the intermediate product of the quantization process. Given any frame of hidden output $\mathbf{h}_{t}$, the integer vector $\mathbf{z}_{t}=[z_{t}^{1},...,z_{t}^{n_{q}}]$ is generated by Residual Vector Quantization (RVQ) (Zeghidour et al., 2021), where $n_{q}$ denotes the number of vector quantization layers. Iteratively, each element $z_{t}^{k}$ is the index among all pre-learned and fixed $k$-th level quantizer vectors $\{\mathbf{q}_{k}^{*}\}$ that has the smallest L2 distance to the residual between $\mathbf{h}_{t}$ and the sum of all previous chosen quantizer vectors $\{\mathbf{q}_{j}^{z_{t}^{j}},j=1,...,k-1\}$. With the discrete representation $\mathbf{z}_{t}$, $\mathbf{\hat{h}}_{t}$ is reconstructed as a close estimation of $\mathbf{h}_{t}$ that can be used to recover $\mathbf{x}_{t}$ with the decoder.
$z_{t}^{k}=\operatorname*{arg\,min}_{m}\text{Distance}(\mathbf{h}_{t}-\sum_{j=1}^{k-1}\mathbf{q}_{j}^{z_{t}^{j}},\mathbf{q}_{k}^{m});\quad\mathbf{\hat{h}}_{t}=\sum_{j=1}^{n_{q}}\mathbf{q}_{j}^{z_{t}^{j}};\quad 1\leq k\leq n_{q}$ (2)
The discrete representation of all audio frames $\mathbf{z}\in\mathbf{Z}^{T\times n_{q}}$ is a matrix and needs to be converted into a sequence before being processed by LM: it is simply flattened as a sequence, in which every $n_{q}$ element for one frame is consecutive. Without specifically stated, we set $n_{q}=3$ in our experiments. As the waveform can be recovered from $\mathbf{z}$ with a neural codec decoder, the rest of this paper mainly discusses how to predict the audio token sequence $\mathbf{z}$ using LLM techniques. As UniAudio intends to generate both speech and non-speech content, we build the codec model on our own and with broader data coverage. Details of our codec configuration is in Appendix E.
2.1.2 Other modalities
Besides audio, other modalities considered in UniAudio also need to be represented as sequences. In addition, most of these sequences are transformed into discrete ones through tokenization. The serialization and tokenization of these input modalities, along with their key features, are briefly summarized as below.
Phoneme: Phonemes are the basic units of speech pronunciation in linguistics. Phoneme sequences have multiple sources: (1) when only text is available, phoneme sequence without duration information can be obtained by text-to-phoneme mapping using a pronunciation dictionary; (2) when only speech is available, phoneme sequence with duration information is obtained by beam search of the DNN-HMM system (Hinton et al., 2012); (3) when both text and speech are available, phoneme sequence with duration information is obtained by forced alignment of the DNN-HMM system  [^2] CMUDict (http://www.speech.cs.cmu.edu/cgi-bin/cmudict) is adopted as the pronunciation dict; kaldi recipe (https://github.com/kaldi-asr/kaldi/tree/master/egs/librispeech/s5/local/chain/run_tdnn.sh) is adopted to build the deep neural network-hidden Markov model (DNN-HMM) system..
MIDI: MIDI (Zhang et al., 2022) is widely used for singing voice synthesis tasks. F0 and duration information are included in the MIDI. We use the duration information to flatten the F0 sequence, so that the frame-level F0 sequence is obtained.
Text: Text acts as a effective carrier of human instructions in audio generation tasks (Yang et al., 2023a; Copet et al., 2023). In this work, these textual instructions are represented as continuous embeddings derived from pre-trained text LLM (Raffel et al., 2020), as these embeddings contain rich textual semantics. Processing these continuous embeddings with LLM is further clarified in Section 2.3333The encoder of T5 (https://github.com/google-research/text-to-text-transfer-transformer) is used to extract the continuous text embeddings..
Semantic Token: The semantic tokens are derived from the continuous embeddings output by audio self-supervised learning (SSL) models. These continuous representations are highly informative and can be adopted in both speech understanding (Rubenstein et al., 2023) and generative tasks (Borsos et al., 2023). Following Huang et al. (2023b), these continuous representations are tokenized by performing K-means clustering (Hsu et al., 2021) over these continuous representations. Since the continuous representations are frame-level, the semantic tokens also encode duration information [^4] The 9-th layer hidden output of Hubert (Hsu et al., 2021) is adopted as the semantic token representations (https://github.com/facebookresearch/fairseq/hubert). The number of clusters for K-means is 500..
2.2 Unified Task Formulation
Table 1: Sequence formats of all tasks supported by UniAudio. Text color represents modality. black: audio; green: phoneme; blue: MIDI; purple: text; brown: semantic token. $\clubsuit$ means tasks that generate audio with deterministic length. $\diamondsuit$: means tasks that are only included in the fine-tuning stage. The speaker prompt is a 3-second speech and is used to represent the speaker identification.
Task Conditions Audio Target
Text-to-Speech (TTS) (Wang et al., 2023a) phoneme, speaker prompt speech
Voice Conversion (VC) ${}^{\clubsuit}$ (Wang et al., 2023e) semantic token, speaker prompt speech
Speech Enhancement (SE) ${}^{\clubsuit}$ (Wang et al., 2023b) noisy speech speech
Target Speech Extraction (TSE) ${}^{\clubsuit}$ (Wang et al., 2018) mixed speech, speaker prompt speech
Singing Voice Synthesis (SVS) (Liu et al., 2022) phoneme (with duration), speaker prompt, MIDI singing
Text-to-Sound (Sound) (Yang et al., 2023c) textual description sounds
Text-to-Music (Music) (Agostinelli et al., 2023) textual description music
Audio Edit (A-Edit) ${}^{\clubsuit\diamondsuit}$ (Wang et al., 2023d) textual description, original sounds sounds
Speech dereverberation (SD) ${}^{\clubsuit\diamondsuit}$ (Wu et al., 2016) reverberant speech speech
Instruct TTS (${}^{\diamondsuit}$ (Guo et al., 2023) phoneme, textual instruction speech
Speech Edit (S-Edit) ${}^{\diamondsuit}$ (Tae et al., 2021) phoneme (with duration), original speech speech
For all tasks considered in UniAudio, the target audio is generated based on given conditions. With the same target modality, i.e., audio, it is the conditions that define different audio generation tasks. However, even with the variance in conditions, all tasks can still be uniformly formulated as sequential modeling tasks that can be processed by LLM: both the target audio and the conditions are first transformed as sub-sequences and spliced as [conditions, target] sequences to be processed.
UniAudio supports 11 audio generation tasks in total. The sequential formats of each task are defined in Table 1, in which the sub-sequences of all modalities are derived as in Section 2.1. However, due to the unique configurations of each task, some of the condition sub-sequences are subject to task-specific pre-processing operations during the tokenization. For audio, these operations are mainly for data corruption, such as adding noise, reverberation, and speech mixed with other speakers in the raw audio before tokenization. For phoneme and semantic tokens, duration information is reserved by default but can also be removed. For singing voice synthesis and speech edit tasks, the duration information of phoneme is used. For TTS and I-TTS tasks, the duration information is not used. For MIDI, the duration information is used repeat the F0 sequence. For text embeddings, no operations are applied in this work.
To avoid ambiguity, some special discrete tokens (enclosed by <>) are inserted to indicate (1) the start and end of the whole sequence; (2) the start and end of each sub-sequence of a certain modality; and (3) the task identifier. For example, for a text-to-sound task sequence that generates target audio based on textual description, the whole sequence is like: <start> <sound_task> <text_start> text_sequence <text_end> <audio_start> audio_sequence <audio_end> <end>.
2.3 Multi-Scale Transformer
Previous work on LLM-based audio generation (Copet et al., 2023) advocates to modeling the discrete audio tokens as flattened sequences. If so, these sequences are processed in the length of $T\times n_{q}$, which is highly challenging considering the quadratic space complexity of Transformer (Vaswani et al., 2017) with respect to the lengths. Inspired by Yu et al. (2023), a multi-scale Transformer architecture is specifically designed for discrete audio sequences, which is a hierarchical model that processes the inter- and intra-frame correlation by global and local Transformer modules separately. An overview of the proposed architecture is in Figure 1. Instead of processing the whole flattened sequence token-by-token like prior works (Kharitonov et al., 2023), the multi-scale transformer considers patches (i.e., every consecutive $n_{q}$ token) as the global modeling units and then handles the tokens within each patch locally. Note that both the global and local Transformers are causal.
For audio token sequences, each patch accounts for $n_{q}$ consecutive audio tokens that exactly represent one audio frame. First, as suggested in Equation 2, regardless of the exact choices of each quantization vector $\mathbf{q}_{*}^{z_{t}^{*}}$, it is the summed quantization vector $\mathbf{\hat{h}}_{t}$ that is used to represent the audio frame. Thus, in the embedding stage, each patch (a.k.a., frame) is represented by the summed vector of the corresponding embeddings before entering the global Transformer. Second, the global Transformer is to predict audio frame-by-frame: to predict the frame $\mathbf{x}_{t}$, it outputs the continuous representations that include frame $\mathbf{x}_{t-1}$ and all previous content. These continuous representations will be further processed by the local Transformer. Third, also as in Equation 2, given the hidden representation $\mathbf{h}_{t}$, the acquisition of $\mathbf{z}_{t}$ is independent of any hidden output other than $\mathbf{h}_{t}$. Inspired by this, it is reasonable to predict the discrete tokens for frame $\mathbf{x}_{t}$, a.k.a., patch $\mathbf{z}_{t}$, only with the hidden output of global Transformer corresponding to frame $\mathbf{x}_{t-1}$. To be more detailed, as the acquisition of each token $z_{t}^{k}$ is auto-regressively dependent on its prior tokens $\{z_{t}^{j}|j<k\}$, a local Transformer is adopted to predict the patch sequence $\mathbf{z}_{t}$ in auto-regressive style. During this process, the corresponding vector output by the global transformer acts as a patch-level context, which is linearly transformed and then added to the embedded results of each token $z_{t}^{k}$.
The proposed multi-scale Transformer architecture is also compatible with discrete and continuous sequences besides audio. For all discrete tokens except audio (phoneme, semantic, MIDI and special tokens), each token has independent semantics and thus should account for one patch. So these discrete tokens repeat for $n_{q}$ times to fill each patch. The continuous text embeddings are also repeated for $n_{q}$ times for the same purpose. Additionally, their embedding process is replaced by a linear transformation while their predicting targets for local Transformer are consecutive special tokens <continuous_token>.
The design of the proposed multi-scale Transformer can effectively reduce computational complexity. First, the equivalent sequence length for the global Transformer is reduced from $T\times n_{q}$ to $T$, which makes the global modeling cost independent to $n_{q}$ and thus the adoption of a larger $n_{q}$ becomes feasible. Second, the intra-patch computation to generate the discrete tokens for each frame is offloaded to the local Transformer. The computation on the local transformer is comparatively light since it only processes the very short sequence (fixed to the length of $n_{q}$) and empirically has fewer parameters than the global Transformer by design.
Figure 1: Overview of UniAudio (left) and multi-scale Transformer architecture (right). <e> represent the end of the sequence. $z_{t}^{k}$ denotes the k-th audio token at t-th frame.

### v6

# AudioSolver: Towards Universal Audio Generation by Large-scale Audio Language Models Pre-training

@@ Abstract
Language models (such as GPT-series) have been shown powerful ability in natural language understanding by large-scale generative pre-training strategy. In this study, we formulate all of the audio generation tasks (such as Text-to-speech (TTS), Voice conversion (VC), Speech separation (SS), Text-to-audio (TTA) and Sing voice generation (SVG)) as a sequence modeling problem. We present AudioSolver, a universal audio language model trained by a multi-task learning manner, which faces the following hinders: (1) a proper agent model that can transfer speech, music, sound, and sing data into one shared latent space. (2) a language model that can deal with long sequences. To solve these problems, we first train a universal audio codec model in large-scale audio-only datasets. Then, we propose a multiscale GPT structure to solve the long sequence problem in audio generation. We scale the training dataset to 150k hours and the model size into 1B. We demonstrate that our universal audio language model has strong generalization ability, which got the SOTA performance in many zero-shot tasks. Furthermore, we find that the pre-trained audio language model can be easy to fine-tune in a new task. To facilitate the research in audio language models, we release AudioSolver, the first open-source audio language model tools, which empowers humans to train audio language models with unprecedented ease.
@@ 1 Introduction
Large language models (LLMs), brown2020language; chowdhery2022palm; OpenAI have been shown powerful language understanding ability, which can solve many natural language processing (NLP) related tasks. Large-scale generative pre-training strategy and increased model parameters are two important factor for the sucess of LLMs. Inspired by the success of LLMs, audio language models got a lot of attention in the context of audio generation. AudioLM borsos2023audiolm is one of the representative work. AudioLM introduced a hierarchical approach which combines two types of audio tokens, with high-level semantic tokens extracted from W2V-BERT chung2021w2v being used to condition the generation of low-level acoustic codes from a neural audio codec model zeghidour2021soundstream. Following AudioLM, a lot of similar works have been proposed to solve audio generation tasks, e.g. VALL-E wang2023neural and SPEARTTS kharitonov2023speak proposed to solve text-to-speech (TTS) task; MusicLM agostinelli2023musiclm and MusicGen copet2023simple proposed to solve text-to-music task; AudioGen kreuk2022audiogen proposed to solve text-to-sound task; Make-A-Voice huang2023make proposed to solve voice conversion tasks. Although these models can get amazing performance for special tasks, their target is far from the LLMs, that is, using a unified model to solve all of tasks. Furthermore, to reserve the acoustic information that influence the user’s auditory sense in the audio, residual vector quantization (RVQ) technique zeghidour2021soundstream is used. RVQ reduces the information loss through increasing the VQ layer, which also increases the burden for generation models (we refer it as long sequence modeling problem). To avoid modelling a very long sequence, VALL-E and AudioLM are both proposed to training a multi-stage models. SPEAR-TTS and MusicGen train a one single stage model, but such models require higher GPU memory. In our practice, training a GPT-like models to deal with long sequence (more than 2000 tokens) is unstable. Thus, long sequence modelling problem is a big hinder in audio language model fields.
Table 1: The audio generation tasks supported by UniAudio and prior works. * means the prior works that do not claim to support that task but we believe they have the potential to. TTS: text-to-speech; VC: voice conversion; SE: speech enhancement; TSE: target speech extraction; SVS: singing voice synthesis; Sound: text-to-sound generation; Music: text-to-music generation; A-Edit: Audio Edit; SD: speech dereverberation; I-TTS: Instructed TTS; S-Edit: Speech edit.
Model TTS VC SE TSE SVS Sound Music A-Edit SD I-TTS S-Edit
YourTTS (yourtts)
VALL-E (valle)
MusicLM (valle)
SPEARTTS (speartts)
NaturalSpeech2 (naturalspeech2)
Make-A-Voice (makeavoice)
Maga-TTS (megatts)
VoiceBox (voicebox)
AudioLDM2 (audioldm2)
SpeechX (speechx)
UniAudio (ours)
In this study, we adopt the idea of large-scale multi-task pre-training from GPT-series radford2018improving; radford2019language; brown2020language. We formulate all of audio related generation tasks as a unified audio language model problem, where we meet with the following research problems: (1) how to choose a proper latent space that different audio data (such as speech, sound, music and sing) can be mapping into the shared discrete space; (2) how to solve the long sequence modeling problem when we want to train a one stage model while keeping more acoustic details; (3) whether large-scale multi-task training can bring improvement; (4) whether pre-trained models can be used for new tasks. In this paper, we will answer these questions. The contributions of this work are:
  • $\bullet$
    We propose a universal audio generation framework to solve all of the audio related generation tasks. To our best knowledge, this is the first work to consider using a language model to solve all of the audio generation tasks.
  • $\bullet$
    To solve the long sequence modelling problem in audio generation, we propose a multiscale GPT structure, that using a global GPT to handle the semantic information modeling, and a local GPT to handle the fine-grained acoustic information modeling. By using our multiscale GPT structure, more acoustic details can be modeling without introducing long sequence.
  • $\bullet$
    Our universal audio language models perform strongly generalization ability. We evaluate our model on zero-shot TTS, zero-shot voice converison, zero-shot speech enhancement, zero-shot speaker extraction, zero-shot sing voice generation, zero-shot text-to-sound and music generation tasks. Experimental results show that our proposed method got the SOTA performance.
  • $\bullet$
    Our pre-trained models can be easy to finue-tune on new tasks. We fine-tune our pre-trained models on speech edit, audio edit, speech dereverberation and instruct-controllable expressive TTS (InstructTTS for short) tasks.
2 AudioSolver
In this section, we introduce the details of AudioSolver. We first introduce how to formulate all of the generation tasks as the sequence modelling problem. Then we give the details of our proposed multi-scale GPT structure to deal with any length sequence. The details of AudioSolver as Figure 1 shows.
2.1 Universal Input Representation
AudioSolver is a universal audio language model that can perceive general modalities, follow instructions, learn in context, and generate outputs. The context can be the phone-sequence, task label, instruction, audio sequence and so on. Based on the previous context, the model learns to generate audio in an auto-regressive manner. Our AudioSolver’s backbone is a Transformer-based causal language model. Following GPT-1, we use the special token to represent different task. We classify the context input into two groups: the first group is discrete sequence, such as phone sequence, audio sequence, task label. The second group is the continuous embedding representations, such as the text representations extracted by a T5 encoder (raffel2020exploring). For these discrete sequence, we directly define a learnable embedding table for them, the embedding table can be optimized with the network. For these continuous embedding, we use a linear layer map them into the same dimension as the learnable embedding dimension, then combine with other embedding sequence. Figure 1 (a) shows six tokenizers to help us realize universal input representation. LabelTokenizer is defined by ourselves, in this study, we define the first 128 tokens in the vocabulary as special tokens, which supports users to define any task label. The later token will start from 128. For PhoneTokenizer and MIDITokenizer, we just follow the map dict that transfer the symbol sequence into token sequence. For SemanticTokenizer, we use the pre-trained HuBERT model and a k-means cluster to quantize the audio signal into semantic token sequence. The HuBERT model’s downsample rate is 320, which means that 1 second audio signal with 16khz sample rate will produce 50 tokens. For AudioTokenzier, we use the neural audio codec models to compress the audio signal into discrete representations. The details of audio codec models will be introduced in the next section.
Figure 1: The overview of AudioSolver.
2.2 Universal Neural Audio Codec
In this study, audio codec models play an important role, which can compress any type audio data (speech, sound, music, sing) into a shared discrete latent space. By using the codec models, we can use the audio language models to generate any type audio. In general, a neural audio codec model consists of three part: (1) an encoder $E$ that takes the audio signal $\bm{x}$ and generates a latent feature representation $\bm{z}$; (2) several residual vector quantization (RVQ) layers that produces a compressed representation $\bm{z}_{q}$; and a decoder that recovers the audio signal $\bm{\hat{x}}$ from the compressed latent representation $\bm{z}_{q}$. In practice, each quantization layer will maintain a codebook (a group of learnable embedding vector). Thus $\bm{z}_{q}$ can be denoted by the index of embedding vector. Assume the down-sample times is $S=320$ in the encoder, and the number of VQ layers $n_{q}=3$, so that 1 second 16khz audio data can be represented by $50\times 3$ tokens. Although Encodec have release the pre-trained models, Encodec results in poor audio quality, which has been verified in VALL-E. HiFi-Codec can bring better audio quality, but it only be trained in speech data.
Table 2: Performance comparison between different universal audio codec models.
Model $n_{q}$ PESQ STOI PESQ STOI PESQ STOI PESQ STOI PESQ STOI
Type - Speech Sound Music Sing AVE
Encodec 8 2.18 0.79 2.23 0.48 1.86 0.57 1.95 0.76 2.05 0.649
I-Encodec 8 3.36 0.878 2.67 0.54 2.31 0.65 3.49 0.89 2.95 0.739
I-Encodec 4 3.11 0.86 2.5 0.51 2.08 0.59 3.27 0.86 2.73 0.71
I-Encodec 3 2.96 0.85 2.42 0.49 1.99 0.57 3.13 0.85 2.62 0.69
HiFiCodec 4 3.21 0.875 2.58 0.53 1.9 0.57 3.5 0.87 2.79 0.71
In this study, we first train a universal audio codec models. We explore two types audio codec models: (1) Improved Encodec model (I-Encodec), we use the same encoder-decoder backbone from open source audio codec model  [^1] https://github.com/facebookresearch/encodec, combined with our improved multi-scale frequency domain discriminator (refer to the Appendix); (2) universal HiFi-Codec model, which we follow the official open source code  [^2] https://github.com/yangdongchao/AcademiCodec, the difference is that we use more audio data. Table 2 shows the reconstruction performance comparison between Encodec and our trained codec models. We can see that our codec models are better than Encodec. Furthermore, using more VQ layers can bring better reconstruction performance. Theoretically, the limit of audio generation model is the reconstruction performance. However, using more VQ layers also bring the burden for the generation models. In the following, unless specifically stated, we follow the setting of SPEARTTS, that using I-Encodec with 3 VQ layers.
Table 3: The comparison between different model structure.
Structure Codec MOS ($\uparrow$) Max audio duration Min GPU memory Training Stage
AudioSolver I-Codec-3 - 30s 16G 1
AudioSolver I-Codec-4 0.64 30s 16G 1
AudioSolver I-Codec-8 0.64 30s 16G 1
VALL-E Encodec-8 - 20s 32G 2
SPEARTTS I-Codec-3 - 10s 64G 1
SPEARTTS I-Codec-8 N/A 1s 64G 1
2.3 Multi-scale GPT Model
Long sequence modeling problem (due to the hierarchical representations in audio codec) is one of the core problems in audio language models. Previous works try to solve this by training multi-stage models, e.g. AudioLM tries to train three separated models, VALL-E trains two independent models. SPEARTTS and MusicGen both find that flattening all codebooks into one sequence brings the best performance. The drawback is that such a pattern will result in long sequences and cost large-scale GPU resources. To combine the advantage of flattening patterns and reducing the sequence length, we propose a multiple-scale GPT structure inspired by yu2023megabyte. Our motivation is that the tokens belonging to different codebooks should enjoy the same semantic information along the same time. Specifically, assuming the audio codec sequence is $a_{1,1},a_{1,2},a_{1,3},...,a_{n,1},a_{n,2},a_{n,3}$ after the flattening operation, where $a_{i,1},a_{i,2},a_{i,3}$ should enjoy the same semantic guidance $s_{i}$. Based on this assumption, we first use a global GPT to model the semantic information (the sequence length is $n$, not previous $3\times n$), then we use a local GPT to generate $a_{i,1},a_{i,2},a_{i,3}$ (the sequence length is 3). By using our model, the max sequence length is $n$, which effectively removes the long sequence problem caused by multi-codebooks. Figure 1 (b) presents the overview of the multi-scale GPT model, which consists of three parts: (1) Patch embedder that combine multi-sequence as one patch (we directly use the concatenate operation along the feature dimension); (2) Global GPT module that using patched features to model the semantic-level information; (3) Local GPT module that using the semantic-level features to model acoustic details. To make the local GPT can parallel deal with all of patch, we repeat other types of tokens $n_{q}$ times, so that the semantic information can be align very well.
2.3.1 Patch Embedder
Assuming the input sequence is $\{x_{i}\}_{i=0}^{i=T-1}$. Patch Embedder aims to reshape the sequence into a patch sequence with the length of $K=\frac{T}{P}$, where $P$ denotes the patch size. Firstly, a global embedding table $E_{g}\in\mathbb{R}^{V\times D_{g}}$ is used to map the input sequence into $\{\bm{x}_{i}\}_{i=0}^{i=T-1}$, where $\bm{x}_{i}\in\mathbb{R}^{D_{g}}$. The positional embedding is used, thus the hidden embedding can be $\bm{h}_{i}^{embed}=\bm{x}_{i}+\bm{p}_{i}$. The Patch Embedder aims to reduce the sequence length of $\bm{H}^{embed}$ by patch concatenate:
$\bm{H}^{embed}\in\mathbf{R}^{T\times D_{g}}\rightarrow\bm{H}^{global-in}\in\mathbb{R}^{K\times(P\cdot D_{g})}$ (1)
2.3.2 Global GPT and Local GPT
After we get the patched features, we can directly feed these features into the global GPT models (a decoder-only Transformer model).
$\bm{H}^{global-out}=GPT^{global}(\bm{H}^{global-in})\in\mathbb{R}^{K\times(P\cdot D_{g})}$ (2)
Then we reshape $\bm{H}^{global-out}$ into the shape of $K\times P\times D_{g}$. Furthermore, we combine $\bm{H}^{global-out}$ with the local embedding (by a local embedding table $E_{l}\in\mathbb{R}^{V\times D_{l}}$):
$\bm{H}^{local-in}=\bm{W}_{g}\bm{H}^{global-out}+\bm{H}^{embed-local}$ (3)
where $\bm{W}_{g}$ denotes the linear mapping layer. Similarly, the local GPT is also a decoder-only Transformer.
$\bm{H}^{local-out}=GPT^{local}(\bm{H}^{local-in})\in\mathbb{R}^{K\times P\cdot D_{l}}$ (4)
Finally, we can compute the probability distribution over the vocabulary at each position based on $\bm{H}^{local-out}$. Lastly, we can calculate the Cross Entropy loss with the ground truth label.